OpenStudy (anonymous):

I have a fraction that has 2/5 and below that is a line and an 8 then there is = 1.2 and a line and N for the missing number and I am so lost

OpenStudy (anonymous):

Can you type the question exactly like it says in the book? Because I can't understand it either.

OpenStudy (heisenberg):

like this? \[\frac{\frac{2}{5}}{8} = \frac{1.2}{N}\] you can try messing with equation editor feature, if you want. all I did was type frac{numerator}{denominator}